Sam earned 1155 in interest on a total investment 15000 this past year, if part of the money was invested at 6% and the remainder at 9% how much was invested at 6%
Answer by Earlsdon(6287)

Let $x = the amount invested at 6%, then the remainder of ($15000 - $x) = the amount invested at 9%. Changing the percents to their decimal equivalents (6% = 0.06 and 9% = 0.09), you can write the equation for this situation:
0.06x + 0.09(15000-x) = 1155 Simplify and solve for x.
0.06x + 1350 - 0.09x = 1155
-0.03x + 1350 = 1155 Subtract 1350 from both sides of the equation.
-0.03x = -195 Divide both sides by -0.03
x = $6500
The amount invested at 6% was $6500